OpenStudy (anonymous):

Solve: 2(x + 4) − (x + 1) = 10

OpenStudy (anonymous):

2(x+4) multiply those

OpenStudy (anonymous):

so thats 2x+8

OpenStudy (anonymous):

Distribute. 2(x + 4) − (x + 1) = 10 2(x + 4) = 2x + 8 -(x + 1) = -x - 1 2x + 8 - x - 1 = 10 Combine like terms. x + 7 = 10 Subtract 7 from both sides. x = 3

OpenStudy (anonymous):

basically what he said ^